CRB Meeting

  • strict warning: Declaration of views_handler_field_user_name::init() should be compatible with views_handler_field_user::init(&$view, $data) in /home/aubreymoore/guaminsects.net/anr/sites/all/modules/views/modules/user/views_handler_field_user_name.inc on line 61.
  • strict warning: Declaration of views_handler_argument_many_to_one::init() should be compatible with views_handler_argument::init(&$view, $options) in /home/aubreymoore/guaminsects.net/anr/sites/all/modules/views/handlers/views_handler_argument_many_to_one.inc on line 169.
CRB MEETING # 10 *08/12/13* *Trappers: R. Cabrera & J. Diego* * Monitored trap route from Afame, Sinajana to Piti. 50 Traps checked, (14) beetles total: (7) old Asan Baseball Field, and (7) Piti Shooting Range. *Sanitation:* *K. Leon Guerrero, K. San Nicolas, & DJ Somol* * Scouted the Leo Palace Resort area. Four snags were cut down but one was identified positive for crb material: (0) Eggs, (0) 1^st , (26) 2^nd , (16) 3^rd , (0) Pupa, and (1F) Beetle. Half day scouting on approximately 20acres. More sanitation work needed in the area. *M. Hara & RJ San Miguel * * Completed sanitation work at Mercy Heights from last Friday’s mission. Materials were taken to Oka Point. More sanitation work needed in the area. *V. Benavente & C. Crisostomo* * Rapid Response: Aaron Palik; negative for infestation of compost around property fence. Feeding damages are visible on property and surrounding area. * Scouted from Perez Villa to Chln. La Chance. Heaving feeding damages seen in certain areas. Scouted approximately 20 acres. May require barrel trap deployment to some areas. Scouting: 40 Sq. Ac. Total *Pesticide: 35 Gallons total* *D. Diego, K. Leon Guerrero, K. San Nicolas, & DJ Somol* * Applied approximately 35 gallons of pesticide to the crowns of thirty-three coconut trees that line the entranceway of the UOG Triton farm in Wusstig. Rd., Yigo. *Others:* *M. Hara & RJ San Miguel: *Assisted R. Quitagua at UOG the first half of the day. K. San Nicolas * *J. Santos: *Contacted forestry department regarding assistance in the controlled burning of Oka Point. Was informed that all forestry personnel are backed up on projects due to a two-week delay at Adelup Compound. Instructed to acquire a burning permit from the Tamuning Fire Station and request for standby while burning for emergency purposes. * *P. Roberto: *Contacted and informed that we are no longer accepting green waste in to the Oka Point compound and that it must be taken to either Primo’s or Limtiaco’s landfill. *TO DO LIST FOR: 8/13/13* *Trappers: *Will monitor traps from Nimitz area heading southbound. *Sanitation: *Scouting and Sanitation in the Merizo village area. *Pesticide:* *D. Diego & K. San Nicolas – *Adelup pesticide application to be finished off early tomorrow morning. Operation will cease at 7:00am.
Groups: